OpenStudy (monsterisenergy):

The volume V of the cylinder is given by the expression V = πR2L . The volume and length of the cylinder are measured as V = 15.0 ± 0.5 cm3 L = 20.0 ± 0.1 cm. Calculate the radius of the cylinder, with its uncertainty. How do you find the UNCERTAINTY?

OpenStudy (monsterisenergy):

And the R is squared.

OpenStudy (monsterisenergy):



OpenStudy (monsterisenergy):

@Mashy @Abhisar @aaronq

OpenStudy (aaronq):

First find the radius cuz you need it \(V=\pi r^2L\rightarrow r=\sqrt{\dfrac{V}{\pi L}}=0.489~cm\) then when you multiply or divide you have to use this formula to find the uncertainty: \(\large \dfrac{\Delta x}{x}=\sqrt{(\dfrac{\Delta y}{y})^2+(\dfrac{\Delta z}{z})^2}\) So for your problem it looks like this: \(\large \dfrac{\Delta r}{r}=\sqrt{(\dfrac{\Delta V}{V})^2+(\dfrac{\Delta L}{L})^2}\) \(\large \dfrac{\Delta r}{0.489}=\sqrt{(\dfrac{0.5}{15.0})^2+(\dfrac{0.1}{20.0})^2}\) \( \large \Delta r =0.017 \approx 0.02\)

OpenStudy (monsterisenergy):

Thanks a lot !!!!!